OpenStudy (anonymous):

Solve for x: 3|x − 3| + 2 = 14

OpenStudy (anonymous):

Is the answer x = −1, x = 7

OpenStudy (anonymous):

14 - 2 = 12 12/3 = 4 x = 7 To check: (7 - 3) = 4 4(3) = 12 + 2 = 14

OpenStudy (anonymous):

Okay, ad the -1 part, I understood the answer was 7

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

3|x − 3| + 2 = 14 3|x − 3| = 12 |x − 3| = 4 now x-3 = 4 OR x-3 = -4

OpenStudy (anonymous):

But 3|x - 3| + 2= 14 would be the equation and -14 can't be involved with absolute value so wouldn't it be no solution?

OpenStudy (anonymous):

I meant with the negatives... Idk, but what about the 7?

OpenStudy (anonymous):

nope the abs|| are the only thing that could be on the other side |x|=|-x|

OpenStudy (anonymous):

Oh, so I'm correct :)

OpenStudy (anonymous):

-1 and 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How many solutions does the equation 2|x − 3| − 8 = −8 have? So with this one, there are zero solutions, right?

OpenStudy (anonymous):

no you have one solution try it

OpenStudy (anonymous):

So problems like these have only one solution? |2x − 3| = −17 right? I just need conformation! :D

OpenStudy (anonymous):

|2x − 3| = −17; no solution

OpenStudy (anonymous):

Okay... the other one had a solution because the absolute value wasn't alone?

OpenStudy (anonymous):

2|x − 3| − 8 = −8 when you add +8 2|x − 3| = 0 ; so the negative goes away

OpenStudy (anonymous):

Oh.... thank you! That's all!

OpenStudy (anonymous):

no problem